OpenStudy (anonymous):

what is the equation of the tangent line to the graph of y = (3e^2x) +1 at the point (0,4)

OpenStudy (anonymous):

y=6x+4

OpenStudy (anonymous):

y=(3e^2x) +1 , y'=6e^2x When x=0, y'=6 <--- this is the slope of the tangent line y=mx+c y=6x+c When x=0, y=4 c=4 The equation of the tangent line is y=6x+4